# The 230 GHz Variability of Numerical Models of Sagittarius A* II. The Physical Origins of the Variability

### v1

@@ 2.2 GRRT
We raytrace the GRMHD simulation snapshots and produce $230$ GHz images using the GRRT code IPOLE (Mościbrodzka & Gammie, 2018). We solve only the unpolarized radiative transfer equation because the $Q$, $U$, and $V$ parameters are out of scope in this study, but they will be included in a subsequent paper. The unpolarized, covariant radiative transfer equation reads
@@
where the subscript $\nu$ refer to the quantity for a given photon frequency, $I_{\nu}$ is the photon intensity, $j_{\nu}$ is the emissivity, and $\alpha_{\nu}$ is the absorptivity. We assume photons travel along the geodesic, where the geodesic equation is given as
@@
Here, $\lambda$ is the affine parameter, and $k^{\nu}$ is the photon four-vector. We use the transfer coefficients in Dexter (2016) to solve Equation 5. The fixed parameters are as follows (Event Horizon Telescope Collaboration et al., 2022): The field of view (FOV) is set to be $200$ $\mu$as. The resolution is $400\times 400$ pixels. The distance from Earth to Sgr A* is taken as $8,178$ parsec, while the mass of Sgr A* is assumed to be $4.154\times 10^{6}\,M_{\odot}$ (Abuter et al., 2019).
@@
Table 1: List of the parameters of interests, recap from Chan et al. (2024).
@@
Table 2: List of black hole parameters of our representative models.
@@ 2.3 Model Selections
The time interval of interest consists of $107$ GRMHD snapshots. Our GRRT spans $10$ $R_{\rm Low}$ and $10$ $M_{\rm Unit}$ for a given $R_{\rm High}$, and we considered $5$ $R_{\rm High}$ and $5$ $\theta$ for a given $a$. The simulation libraries span $5$ different spins, thus we have $\sim$ 12,500$\times$107 = 1,337,500 GRRT images. It is almost impractical to analyze every single one of the images. Instead, we choose representative models that could help us infer qualitatively the parameter dependence of $M_{\Delta T}$ on $R_{\rm Low}$. The selection procedure is as follows: Given a $R_{\rm High}$, $\theta$, and $a$, we first identify if the sequences of models contain a local minimum in $M_{\Delta T}$. If it does, we select $4$ representative models along the sequence — One at $R_{\rm Low}=1$, one at $\text{MIN}[60,R_{\rm High}]$, one at the local minimum, and one in between $\text{MIN}[60,R_{\rm High}]$ and the local minimum. If, however, the sequence of models does not contain a local minimum, we will sample $4$ models evenly from $R_{\rm Low}=1$ to $\text{MIN}[60,R_{\rm High}]$. We summarize the black hole parameters of our representative models in Table 2 for reference.
@@ 3.1 Positive Spins
In Chan et al. (2024), we find that $M_{\Delta T}$ could be reduce by varying $R_{\rm Low}$ for almost all black holes with $a>0$, and that a local minimum exists between $R_{\rm Low}=1$ and $\text{MIN}[60,R_{\rm High}]$. To understand this, we show the images and $230$ GHz light curves of the black hole with $a=+0.94$, $\theta=0.1$ Deg, and $R_{\rm High}=180$ in Figure 1. Model with $R_{\rm Low}=1$ has $M_{\Delta T}=0.147$, while the model with the lowest $M_{\Delta T}$ has $R_{\rm Low}=14.58$ and $M_{\Delta T}=0.04$. At a higher $R_{\rm Low}$, $M_{\Delta T}$ increases from the local minimum to a value of $0.180$, which is even larger than that at $R_{\rm Low}=1$. This is consistent with the ‘first decrease than increase’ trend we found in Chan et al. (2024). The light curves could be easily separated into different domains in time where the $R_{\rm Low}=1$ model is more/less variable than models with higher $R_{\rm Low}$. The model with $R_{\rm Low}=14.58$ has the flattest light curve.
@@
Why does the $R_{\rm Low}=14.58$ model have the lowest variability? And where does the large variability of the higher $R_{\rm Low}$ models originate? We select GRRT snapshots at $t=29,840$$GM/c^{3}$ and show them on the first row of Figure 1. They are marked as squares on the light curves and correspond to the moment where the $R_{\rm Low}=1$ model is more variable than the rest. The major contribution to the $230$ GHz flux is from the photon ring, consistent with our findings in Chan et al. (2024). As $R_{\rm Low}$ increases, the optical depth of the fluid also increases and helps obscure the photon rings, reducing its contribution to the variability. We also note that the higher $R_{\rm Low}$ models at this particular moment are dimmer than the averaged $230$ GHz flux. We illustrate the increase in optical depth with increasing $R_{\rm Low}$ by showing the time-averaged optical depth of the black hole in Figure 2. Here, the scale $1-2$ is mapped, while that from $0-1$ is physical. The $R_{\rm Low}=1$ model has an optically thin photon ring. The photon ring for the $R_{\rm Low}=14.58$ model is marginally optically thin, while those having larger $R_{\rm Low}$ are optically thick. These combined results suggest that the reduced variability is an optical depth effect.
@@
We show in the second row of Figure 1 the images at $t=29,980$$GM/c^{3}$, where the higher $R_{\rm Low}$ models are more variable than the $R_{\rm Low}=1$ and $14.58$ model. These are marked as triangles on the light curves. In the image plane, we find that the higher $R_{\rm Low}$ models show bright transients near the black hole, and they become more visible as $R_{\rm Low}$ is increased. The origins of this transient could be understood through analyzing GRMHD snapshots. We plot the plasma entropy $\mathcal{S}=\text{log}\{P/[\rho^{\gamma}(\gamma-1)]\}$ contours at this particular moment in Figure 3. We find a patch of plasma close to the black hole to have an unusually large $\mathcal{S}$ compared to its surroundings, and it has a shape consistent with the bright transients shown in Figure 3. The $x-z$ slice of the snapshot shows a tube-like structure for this patch of plasma. This is probably a MAD flux eruption caused by reconnections (Dexter et al., 2020b; Scepi et al., 2022, 2024).
@@
Black holes with $a=+0.5$ exhibit similar trends in the $R_{\rm Low}$ vs $M_{\Delta T}$ curves as those of $a=+0.94$. To understand this, we show the $230$ GHz light curves and images for black holes with $a=+0.5$, $\theta=45$ Deg, and $R_{\rm High}=120$ in Figure 4. Like Figure 1, we can divide the light curves into domains where the $R_{\rm Low}=1$ models are more/less variable than the remaining. The model with $R_{\rm Low}=23.68$ is the least variable with $M_{\Delta T}=0.103$, and its light curve is the flattest. We show in the first row, the images at $t=29780$$GM/c^{3}$ (marked as triangles on the light curves) where the $R_{\rm Low}=1$ model is more variable than the remaining. We find, again, that the major contribution to the brightness for the $R_{\rm Low}=1$ model is from the photon ring and that optically thick fluid blocks the photon ring from illuminating for models with a higher $R_{\rm Low}$.
@@
At the beginning of this particular time chunk, the higher $R_{\rm Low}$ models are much more variable. We show in the second row of Figure 4 the images at this instant. We find a prominent, tube-like feature in the images, which becomes more visible as $R_{\rm Low}$ increases further. This is probably a large-scale MHD flux tube created by the MAD flux eruptions — After the black hole saturates in horizon-penetrating magnetic flux, part of the magnetic flux is dissipated via a magnetic reconnection that generates a strong vertical magnetic field, which then pushes back the accretion flow and creates the flux tube (Vos et al., 2023). We illustrate the image of the flux tube viewing at different $\theta$ in Figure 5. Note that $R_{\rm Low}\neq 1$ means the image is slightly optically thicker than the $R_{\rm Low}=1$ model. To our knowledge, this is the first illustration of a MAD eruption flux tube in the image domain. In Figure 6, we show the density contour with magnetic field lines at the same instant. A patch of tube-like, low-density plasma threaded with strong vertical magnetic fields is visible, thus confirming our claim that the tube-like structure is a MAD eruption flux tube.
@@
In short, for black holes with $a>0$, the major contribution to the variability of the $R_{\rm Low}=1$ model is from the photon ring. Increasing $R_{\rm Low}$ increases the optical depth of the fluid, which helps block the photon ring from illuminating. However, it comes with the cost of making MAD flux eruptions more visible, thus increasing $M_{\Delta T}$. The sweet spot occurs when the contributions to $M_{\Delta T}$ from the photon rings and the flux eruptions are combined at the minimal.
@@ 3.2 Negative Spins
In Chan et al. (2024), we observed an increasing trend in $M_{\Delta T}$ when we attempted to increase $R_{\rm Low}$ for almost all black holes with $a<0$. To understand this, we show the black holes with $a=-0.5$, $\theta=0.1$ Deg, and $R_{\rm High}=90$ and show the resulting $230$ GHz light curves and images in Figure 7. We show only the moment where the higher $R_{\rm Low}$ models are more variable. We find that the major contributions to the variability for the $R_{\rm Low}=1$ models are from the photon ring, consistent with our findings in Chan et al. (2024). Increasing $R_{\rm Low}$ also makes the fluid optically thicker, as in the $a>0$ cases, but the fluid contains voids that cannot cover the entire photon ring. We also show the black holes with the same spin but viewed at $\theta=90$ Deg in Figure 8. We see the same image morphology where optically thick gas cannot cover the photon ring. Thus, this phenomenon should be generic across varying $\theta$. The origin of these voids could be easily traced through GRMHD snapshots. We show the density (plasma $\beta^{-1}$) contours in the upper (lower) panel of Figure 9. We observe patches of low-density plasma surrounding the black hole. The plasma $\beta^{-1}$ contour also shows that these voids are highly magnetized, suggesting they might carry considerable magnetic flux.
As in the $a>0$ cases, MAD flux eruptions are more visible as one increases $R_{\rm Low}$, which could be seen as bright strips in the increasing $R_{\rm Low}$ models in Figure 7. We verify this by showing the entropy $\mathcal{S}$ contour in Figure 10. We find a strip of unusually high $\mathcal{S}$ plasma floating away from the black hole, which has a shape consistent with the bright strips shown in Figure 7. We note that the flux eruptions span a larger spatial scale, making them visible even when $R_{\rm Low}=1$. This differs from the $a>0$ cases, where the flux eruptions become visible only when $R_{\rm Low}>1$ and their spatial scale is much smaller.
There is an exception for the above discussion: the model with $a=-0.94$ and $\theta=179.9$ Deg. We find a decreasing trend in the $R_{\rm Low}$ vs $M_{\Delta T}$ curves regardless of the $R_{\rm High}$ assumed. To understand this, we show the images and $230$ GHz light curves of black holes for this set of parameters but with $R_{\rm High}=180$ in Figure 11. The photon rings are still visible for models with $R_{\rm Low}>1$. The flux eruptions are visible even at $R_{\rm Low}=1$. The combined variability due to flux eruptions and photon rings makes the $R_{\rm Low}=1$ model stand out compared to those with $R_{\rm Low}>1$. However, since the increasing trend is quite generic for all of the $a=-0.5$ models and for the $a=-0.94$ models viewed from another angle, we believe that this is a particular case that would not affect our qualitative understanding of the parameter dependence of $M_{\Delta T}$ on $R_{\rm Low}$ for black holes with $a<0$.
@@ 3.3 Zero Spin
We find that $M_{\Delta T}$ are not sensitive to the variation of $R_{\rm Low}$ for black holes with $a=0$, irrespective of the $R_{\rm High}$ assumed. Here, we select the black holes with $a=0$, $\theta=135$ Deg, and $R_{\rm High}=50$ and show the resulting $230$ GHz light curves and images in Figure 12. The optical depth argument still applies, in which the photon rings are blocked by optically thick gas. But the $230$ GHz light curves for models with larger $R_{\rm Low}$ are similar, even though they differ significantly from that of $R_{\rm Low}=1$. The light curves thus reveal why $M_{\Delta T}$ are insensitive to $R_{\rm Low}$. The reason for this behaviour would probably be either the interval of snapshots we selected is when the black holes are at their quiescent state or the black holes with $a=0$ are intrinsically quiet. Nonetheless, given that the current EHT constraint on Sgr A* prefers a black hole with a non-zero spin, the $a=0$ case is seemingly out of interest and would not affect our physical interpretation of the variability of modelling Sgr A*. However, the physics of such insensitivity would be interesting, and we would leave this work for a subsequent paper.

### v2

@@ 2.2 GRRT
We raytrace the GRMHD simulation snapshots and produce $230$ GHz images using the GRRT code IPOLE (Mościbrodzka & Gammie, 2018). We solve only the unpolarized radiative transfer equation because the $Q$, $U$, and $V$ parameters are out of scope in this study, but they will be included in a subsequent paper. The unpolarized, covariant radiative transfer equation reads:
@@
where the subscript $\nu$ refers to the quantity for a given photon frequency, $I_{\nu}$ is the photon intensity, $j_{\nu}$ is the emissivity, and $\alpha_{\nu}$ is the absorptivity. We assume photons travel along the geodesic, where the geodesic equation is given as:
@@
here, $\lambda$ is the affine parameter, and $k^{\nu}$ is the photon four-vector. We use the thermal synchrotron transfer coefficients in Dexter (2016) to solve Equation 6. The fixed parameters are as follows (Event Horizon Telescope Collaboration et al., 2022): The field of view (FOV) is set to be $200$ $\mu$as. The resolution is $400\times 400$ pixels. The distance from Earth to Sgr A* is taken as $8,178$ parsecs, while the mass of Sgr A* is assumed to be $4.154\times 10^{6}\,M_{\odot}$ (GRAVITY Collaboration et al., 2019). Note that emission from plasma with $\sigma>1$ are excluded because of artificial mass injection due to density floors in the simulations.
@@
Table 1: List of the parameters of interest, recapped from Chan et al. (2024).
@@
Table 2: List of black hole parameters for our representative models.
@@ 2.3 Model Selections
The time interval of interest consists of $107$ GRMHD snapshots. Our GRRT spans $10$ $R_{\rm Low}$ and $10$ $M_{\rm Unit}$ for a given $R_{\rm High}$, and we considered $5$ $R_{\rm High}$ and $5$ $\theta$ for a given $a$. The simulation libraries span $5$ different spins, thus we have approximately $12,500$$\times$ $107$ = $1,337,500$ GRRT images. It is almost impractical to analyze every single one of the images. Instead, we choose representative models that could help us infer qualitatively the parameter dependence of $M_{\Delta T}$ on $R_{\rm Low}$. We summarize the black hole parameters of our representative models in Table 2 for reference.
The selection procedure is as follows: Given a $R_{\rm High}$, $\theta$, and $a$, we first identify if the sequence of models contains a local minimum in $M_{\Delta T}$. If it does (e.g., upper left and upper center of Figure 1), we select 4 representative models along the sequence—one at $R_{\rm Low}=1$, one at $\text{MIN}[60,R_{\rm High}]$, one at the local minimum, and one in between $\text{MIN}[60,R_{\rm High}]$ and the local minimum. If, however, the sequence of models does not contain a local minimum (e.g., lower center and lower right of Figure 1), we will sample 4 models evenly from $R_{\rm Low}=1$ to $\text{MIN}[60,R_{\rm High}]$.
@@ 3.1 Positive Spins
In Chan et al. (2024), we find that $M_{\Delta T}$ could be reduced by varying $R_{\rm Low}$ for almost all black holes with $a>0$, and that a local minimum exists between $R_{\rm Low}=1$ and $\text{MIN}[60,R_{\rm High}]$. To understand this, we show the images and $230$ GHz light curves of the black hole with $a=+0.94$, $\theta=0.1$ Deg, and $R_{\rm High}=180$ in Figure 2. We convert images from units of intensity to brightness temperature via:
$T_{b}=\frac{I_{\nu}c^{2}}{2k\nu^{2}}$ (9)
where $I_{\nu}$ is the intensity of each pixel, $\nu=230$ GHz, and $k$ is the Boltzmann constant. The model with $R_{\rm Low}=1$ has $M_{\Delta T}=0.147$, while the model with the lowest $M_{\Delta T}$ has $R_{\rm Low}=14.58$ and $M_{\Delta T}=0.04$. At a higher $R_{\rm Low}$, $M_{\Delta T}$ increases from the local minimum to a value of $0.180$, which is even larger than that at $R_{\rm Low}=1$. This is consistent with the ‘first decrease then increase’ trend we found in Chan et al. (2024). The light curves could be easily separated into different domains in time where the $R_{\rm Low}=1$ model is more/less variable than models with higher $R_{\rm Low}$. The model with $R_{\rm Low}=14.58$ has the flattest light curve.
@@
Why does the $R_{\rm Low}=14.58$ model have the lowest variability? And where does the large variability of the higher $R_{\rm Low}$ models originate? We select GRRT snapshots at $t=29,840$$GM/c^{3}$ and show them on the first row of Figure 2. They are marked as squares on the light curves and correspond to the moment where the $R_{\rm Low}=1$ model is more variable than the rest. The major contribution to the $230$ GHz flux is from the photon ring, consistent with our findings in Chan et al. (2024). As $R_{\rm Low}$ increases, the optical depth of the fluid also increases and helps obscure the photon rings, reducing its contribution to the variability. We also note that the higher $R_{\rm Low}$ models at this particular moment are dimmer than the averaged $230$ GHz flux. We illustrate the increase in optical depth with increasing $R_{\rm Low}$ by showing the time-averaged optical depth of the black hole in Figure 3. The frequency-dependent optical depth along each ray that follows the geodesic is given as an integral:
[…]
where $\lambda_{0}$ ($\lambda_{1}$) is the point of emission (recipient). Note that the absorption coefficient $\alpha_{\nu}$ is evaluated in the fluid rest frame. We compute optical depth images for each snapshot and then take the time average. Here, the scale $1-2$ is mapped, while that from $0-1$ is physical. The $R_{\rm Low}=1$ model has an optically thin photon ring. The photon ring for the $R_{\rm Low}=14.58$ model is marginally optically thin, while those with larger $R_{\rm Low}$ are optically thick. These combined results suggest that the reduced variability is an optical depth effect – the gravitationally lensed photons, even though traveling through the same physical length, sample a larger photon path length due to increased fluid density.
@@
We show in the second row of Figure 2 the images at $t=29,980$$GM/c^{3}$, where the higher $R_{\rm Low}$ models are more variable than the $R_{\rm Low}=1$ and $14.58$ models. These are marked as triangles on the light curves. In the image plane, we find that the higher $R_{\rm Low}$ models show bright transients near the black hole, which become more visible as $R_{\rm Low}$ increases. The origins of this transient can be understood through analyzing GRMHD snapshots. We plot the plasma entropy $\mathcal{S}=\text{log}\{P/[\rho^{\gamma}(\gamma-1)]\}$ contours at this particular moment in Figure 4. We find a patch of plasma close to the black hole that has an unusually large $\mathcal{S}$ compared to its surroundings, and it has a shape consistent with the bright transients shown in Figure 4. The $x-z$ slice of the snapshot shows a tube-like structure for this patch of plasma. This is probably a MAD flux eruption caused by reconnections (Dexter et al., 2020b; Scepi et al., 2022, 2024).
@@
Black holes with $a=+0.5$ exhibit similar trends in the $R_{\rm Low}$ vs. $M_{\Delta T}$ curves as those with $a=+0.94$. To understand this, we show the $230$ GHz light curves and images for black holes with $a=+0.5$, $\theta=45$ Deg, and $R_{\rm High}=120$ in Figure 5. Like Figure 2, we can divide the light curves into domains where the $R_{\rm Low}=1$ models are more or less variable than the remaining ones. The model with $R_{\rm Low}=23.68$ is the least variable with $M_{\Delta T}=0.103$, and its light curve is the flattest.
We show the images in the first row at $t=29,780$$GM/c^{3}$ (marked as triangles on the light curves) where the $R_{\rm Low}=1$ model is more variable than the others. We find, again, that the major contribution to the brightness for the $R_{\rm Low}=1$ model is from the photon ring and that optically thick fluid blocks the photon ring from illuminating for models with a higher $R_{\rm Low}$.
@@
At the beginning of this particular time chunk, the higher $R_{\rm Low}$ models are much more variable. We show in the second row of Figure 5 the images at this instant. We find a prominent, tube-like feature in the images, which becomes more visible as $R_{\rm Low}$ increases further. This is probably a large-scale MHD flux tube created by the MAD flux eruptions — after the black hole saturates in horizon-penetrating magnetic flux, part of the magnetic flux is dissipated via a magnetic reconnection that generates a strong vertical magnetic field, which then pushes back the accretion flow and creates the flux tube (Vos et al., 2024). We illustrate the image of the flux tube viewed at different $\theta$ in Figure 6. Note that $R_{\rm Low}\neq 1$ means the image is slightly optically thicker than the $R_{\rm Low}=1$ model. In Figure 7, we show the density contour with magnetic field lines at the same instant. A patch of tube-like, low-density plasma threaded with strong vertical magnetic fields is visible, thus confirming our claim that the tube-like structure is a MAD eruption flux tube.
@@
In short, for black holes with $a>0$, the major contribution to the variability of the $R_{\rm Low}=1$ model is from the photon ring. Increasing $R_{\rm Low}$ increases the optical depth of the fluid, which helps block the photon ring from illuminating. However, this comes with the cost of making MAD flux eruptions more visible, thus increasing $M_{\Delta T}$. The sweet spot occurs when the contributions to $M_{\Delta T}$ from the photon rings and the flux eruptions are combined at the minimal.
@@ 3.2 Negative Spins
In Chan et al. (2024), we observed an increasing trend in $M_{\Delta T}$ when we attempted to increase $R_{\rm Low}$ for almost all black holes with $a<0$. To understand this, we show the $230$ GHz light curves and images in Figure 8. Here, the black hole parameters are $a=-0.5$, $\theta=0.1$ Deg, and $R_{\rm High}=90$. We show only the moment where the higher $R_{\rm Low}$ models are more variable. We find that the major contributions to the variability for the $R_{\rm Low}=1$ models are from the photon ring, consistent with our findings in Chan et al. (2024). Increasing $R_{\rm Low}$ also makes the fluid optically thicker, as in the $a>0$ cases, but the fluid contains voids that cannot cover the entire photon ring. We also show the black holes with the same spin but viewed at $\theta=90$ Deg in Figure 9. We see the same image morphology where optically thick gas cannot cover the photon ring. Thus, this phenomenon should be generic across varying $\theta$. The origin of these voids could be easily traced through GRMHD snapshots. We show the $x-y$ ($x-z$) density contours in the upper (lower) panel of Figure 10. We observe a flux eruption events as in the $a>0$ case, and we also find patches of low-density plasma surrounding the black hole.
As in the $a>0$ cases, MAD flux eruptions are more visible as one increases $R_{\rm Low}$, which can be seen as bright strips in the increasing $R_{\rm Low}$ models in Figure 8. We verify this by showing the entropy $\mathcal{S}$ contour in Figure 11. We find a strip of unusually high $\mathcal{S}$ plasma floating away from the black hole, which has a shape consistent with the bright strips shown in Figure 8. We note that the flux eruptions span a larger spatial scale, making them visible even when $R_{\rm Low}=1$. This differs from the $a>0$ cases, where the flux eruptions become visible only when $R_{\rm Low}>1$ and their spatial scale is much smaller.
There is an exception to the above discussion: the model with $a=-0.94$ and $\theta=179.9$ Deg. We find a decreasing trend in the $R_{\rm Low}$ vs. $M_{\Delta T}$ curves regardless of the $R_{\rm High}$ assumed. To understand this, we show the images and $230$ GHz light curves of black holes for this set of parameters but with $R_{\rm High}=180$ in Figure 12. The photon rings are still visible for models with $R_{\rm Low}>1$. The flux eruptions are visible even at $R_{\rm Low}=1$. The combined variability due to flux eruptions and photon rings makes the $R_{\rm Low}=1$ model stand out compared to those with $R_{\rm Low}>1$. However, since the increasing trend is quite generic for all of the $a=-0.5$ models and for the $a=-0.94$ models viewed from another angle, we believe that this is a particular case that would not affect our qualitative understanding of the parameter dependence of $M_{\Delta T}$ on $R_{\rm Low}$ for black holes with $a<0$.
@@ 3.3 Zero Spin
We find that $M_{\Delta T}$ is not sensitive to the variation of $R_{\rm Low}$ for black holes with $a=0$, irrespective of the $R_{\rm High}$ assumed. Here, we select the black holes with $a=0$, $\theta=135$ Deg, and $R_{\rm High}=50$ and show the resulting $230$ GHz light curves and images in Figure 13. The optical depth argument still applies, in which the photon rings are blocked by optically thick gas. However, the $230$ GHz light curves for models with larger $R_{\rm Low}$ are similar, even though they differ significantly from that of $R_{\rm Low}=1$. The light curves thus reveal why $M_{\Delta T}$ is insensitive to $R_{\rm Low}$. The reason for this behavior might be that the interval of snapshots we selected is when the black holes are at their quiescent state, or the black holes with $a=0$ are intrinsically quiet. Nonetheless, given that the current EHT constraint on Sgr A* prefers a black hole with a non-zero spin, the $a=0$ case is seemingly out of interest and would not affect our physical interpretation of the variability of modeling Sgr A*. However, the physics of such insensitivity would be interesting, and we will leave this work for a subsequent paper.